OpenStudy (samsan9):

what is the velocity acquired by a free falling object 5 s after being dropped from a rest position? what is the velocity 6 s after?

OpenStudy (midhun.madhu1987):

Here t = 5 s , u = o m/s, g = 9.8m/s^2 v = u + gt Find v from this equation...

OpenStudy (samsan9):

49 m/s?

OpenStudy (midhun.madhu1987):

yes.. :)

OpenStudy (samsan9):

and 6?

OpenStudy (midhun.madhu1987):

just make a try

OpenStudy (samsan9):

58.8?